# AstroLoc: Robust Space to Ground Image Localizer

### v1

@@ 4.2 Query-Satellite Pairwise Loss
Given the cross-domain nature of the problem, we aim to take matching images from the different distributions and applying contrastive learning to encode their relationship in the feature space. To this end, we create a batch $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{B}\}$ of query-database pairs $p_{i}=(q_{i},d_{i})\in(\mathcal{Q},\mathcal{D})$ that have an intersection over union (IoU) higher than a threshold $t_{iou}$ [^6] For the sake of readability, we abused the notation by implying that a matching query and database images from the sets $\mathcal{Q}$ and $\mathcal{D}$ have the same index as the index of the pair they belong to. This trick to simplify the notation is equivalent to have preliminarily sorted the sets $\mathcal{Q}$ and $\mathcal{D}$. (see examples in Fig. 4 top). We ensure that within a batch there are no two pairs that have geographic overlap, so that for a pair $p_{i}\in\mathcal{P}$ all the images from other pairs $p_{j}\in\mathcal{P}\backslash\{p_{i}\}$ are negative examples.
@@ 4.3 Unsupervised Mining
Figure 5: Weighted clusters from unsupervised mining.  The spatial distribution of images from 50 mined clusters, colored by cluster sampling weight. Higher weighted areas (yellow) are sampled more frequently than lower weighted areas (blue).
To fully exploit the much larger corpus of satellite image, we use a second loss inspired by EarthLoc [5], where quadruplets of satellite images are used with a multi-similarity loss [35] to learn robust feature representations. Each quadruplet contains four images of the same place, i.e. positives toward each other, and negatives w.r.t. the images from other quadruplets. To implement this idea, EarthLoc [5] first uses an offline clustering strategy, to divide the database images into visually similar subsets (e.g., a cluster with images of mountains, one with deserts, one with coastlines, etc.). Formally, it creates a set of $K$ clusters $\mathcal{C}_{1},\ldots,\mathcal{C}_{K}$ such that
@@
The clusters are computed with a k-means algorithm using features computed with the model while it’s training, as common practice with mining techniques [2]. Then, at each iteration a batch of quadruplets from one single cluster is created, where the chosen cluster is uniformly sampled among all possible clusters. Formally, we denote this sampled batch of quadruplets as
$\mathcal{H}_{k}=\{h_{1},\ldots,h_{H}\}$ (7)
where the generic $h_{i}=(h_{i_{1}},\ldots,h_{i_{4}})\in(\mathcal{C}_{k},\mathcal{C}_{k},\mathcal{C}_{k},\mathcal{C}_{k})$ is a quadruplet of images from the same location, and the cluster is chosen according to uniform sampling
$k\sim U(1,K)$ (8)
EarthLoc demonstrates that this random sampling of clusters achieves significant improvement over sampling quadruplets from the entire database $\mathcal{D}$. This is because clustering ensures that all the images in the batch, including the ones used as negatives, are visually similar, thus yielding a better signal for a contrastive loss (i.e. that the negatives are harder negatives). However, we argue that the uniform sampling in Eq. 8 leads the model to learn on a data distribution that is geographically very dissimilar from how the photos from astronauts are distributed. Generally, astronauts take few photos of feature-less landscapes like deserts and ice sheets, while often photographing scenes relevant to natural disasters and climate change like glaciers, coastlines and islands.
Figure 6: t-SNE plot of features with and without our mining. Blue dots represent query set features, and other dots represent features of the training quadruplets (made of database images) with and without our mining, (right and left, respectively). Although neither is a perfect blend, our mining allows sampling from a distribution much more intertwined with the query distribution.
To avoid this distribution shift, we propose sampling the cluster to be used to create the batch $\mathcal{H}_{k}$ according to the distribution of the query set $\mathcal{Q}$. We first compute the features from our training queries, assign them to clusters, and then apply a weighted sampling of the clusters weighted by how many queries belong to it. Note that the clusters are computed only on the database, and the queries are simply assigned to them to determine the clusters’ weighting. For example, if there are three clusters A, B and C (see Fig. 4), and 10% of queries are assigned to cluster A, 90% to cluster B, and 0 to cluster C, then the training batches will be sampled 10% from cluster A and 90% from cluster B (see Fig. 4). Formally, let’s denote with $b_{1},\ldots,b_{K}$ the size of the bins derived from assigning the features of the training queries to the $K$ clusters. The sampling of the cluster is then replaced by
@@
Unlike what commonly done in contrastive learning [22, 2, 1, 21], this mining is unsupervised, as the queries’ labels are not used, meaning that we could potentially use all of the 5M existing astronaut photos. In practice we use only training queries to avoid seeing test queries at train time. The second loss used in our pipeline is thus
@@
where $k$ is sampled according to Eq. 9, $\alpha_{2}$ and $\beta_{2}$ are positive hyperparameters, and MUM stands for Multi-similarity with Unsupervised Mining.
To summarize, our unsupervised mining relies on the following steps:
@@
    extract features from database (i.e. satellite imagery);
@@
    compute k-means clustering on these features;
  3. 3.
    extract features from training queries (i.e. astronaut photos), assign them to the clusters;
  4. 4.
    sample database images in proportion to the number of training queries assigned to that cluster.
These steps are repeated periodically every 5000 iterations, increasing training time by less than 10%. This is a data-driven unsupervised mining, where the goal is to sample database images using the training queries distribution (defined in eq. 9). The practical effect of this mining is visible in Fig. 5, where yellower areas represent more informative clusters that are therefore sampled more often.
@@ 5.1 Implementation Details
Training We set the hyperparameters as follows: $t_{iou}=0.2,\alpha_{1}=1,\alpha_{2}=1,\beta_{1}=50,\beta_{2}=50,\lambda_{1}=1,\lambda_{2}=1,K=50$, batch size $=$ 48 (48 pairs for pair loss, 48 quadruplets for MUM loss), learning rate $=$ 5e-5, Adam optimizer. We use as architecture a DINO-v2-base backbone [25] with SALAD [14] and a linear layer to reduce feature dimensionality (from 8448 to 2048) (i.e. a similar model to AnyLoc [17] while being over 10 times lighter as they use DINO-v2-giant). Training runs for 30k iterations. As in [5], the Texas dataset is used as validation.
Evaluation We follow common procedure [5] by doing image retrieval with an augmented dataset (i.e. applying four 90° rotations to each image), and using as metric the recall@N, as the percentage of queries where at least one of the top-N predictions is a correct match to the query.
@@ 5.2 Results
Results clearly show that, whereas previously there was no dominant method between EarthLoc and AnyLoc, the introduction of AstroLoc provides a new state-of-the-art model which significantly outperforms all previous, achieving a near perfect (above 99%) recall@100 on all six evaluation sets. Note that the recall@100 is more relevant in the real world than recall@1, since it is common practice to re-rank the top-N predictions with image matching methods [4, 32]. Some qualitative results of queries and their predictions are in Fig. 1, with many more in the supplementary.
@@ 5.3 Other Space to Ground Use Cases
Table 4: Results of Lost-in-Space satellite problem. Using the dataset from VINSat [20] as queries, we show AstroLoc can also achieve good results on this additional task. AstroLoc-tiny is a tiny version more suitable for deployment on a nanosatellite, with 27M model parameters and 512 dimensional features.
@@
Historical space imagery localization Historical imagery of Earth from space is a unique source of data to understand how Earth has changed over decadal time spans. We perform experiments on 704 queries from early Space Shuttle days (1981-1984), on a worldwide database, and report results in Tab. 5. We empirically find that AstroLoc achieves strong results even on these old photographs, showcasing its robustness to various types of domain changes.
@@ 5.4 Ablations
Training setup We compute ablations on the losses and mining in Tab. 6, where results clearly show the strong impact from each module in our pipeline. Most importantly, the combination of these components presents the optimal characteristic of orthogonality, in that their mixture has notably higher results than each of the singular elements. Features visualizations of the model trained with and without our mining is also shown in Fig. 6.
Method Texas-L Alps-L California-L R@1 R@10 R@100 R@1 R@10 R@100 R@1 R@10 R@100 Pair Loss 83.6 93.1 97.5 87.2 95.0 98.2 85.8 93.7 97.8 MS loss (default mining) 72.4 83.3 91.5 79.4 89.1 94.7 75.4 85.5 92.9 MUM loss 82.2 91.1 97.1 86.9 94.5 97.9 84.1 93.0 97.8 Pair Loss + MS loss 86.5 94.1 97.7 91.5 96.5 98.6 87.8 94.6 97.9 AstroLoc: Pair Loss + MUM loss 91.1 96.2 98.5 94.6 97.8 99.2 92.1 96.7 98.7
Table 6: Ablations on training setup. We show results using different loss formulations (only pair/MS/MUM, and combinations of these) and then the final complete architecture, AstroLoc, comprising the pair loss and MUM loss.
Feature dimensionality Given that feature dimensionality plays a crucial role in any large-scale retrieval task, as it is linearly correlated with the test-time memory footprint and matching time (i.e. the time it takes to process the kNN), we present a plot (Fig. 7) of how results vary when changing the dimension, as we compute results with two versions of the DINO-v2 backbone, base and small. Interestingly, we find that results increase up to the highest dimensionality (i.e. 8448, equal to the input dimension to the final linear layer), showing no plateau at lower dimensions.
Figure 7: Ablation on feature size. Backbone (DINO-v2 small/base) with different feature size/dimension. Stars are models selected for our experiments, blue is AstroLoc-tiny (Sec. 5.3) and orange is AstroLoc in all other tables. R@1 computed on Texas-L.
@@ 6 Conclusions
We tackle the task of Astronaut Photography Localization (APL), obtaining a robust APL model by leveraging a huge set of weakly annotated astronaut photographs, for which we first computed full localization labels. We propose two ways to enhance our model through this data, first adopting them (paired with their satellite counterparts) in an ad-hoc pairwise loss, and second using them to mine ideal training data to pass to our model. We thus obtain a new model, AstroLoc, which shows impressive results on all APL datasets; furthermore, we find that AstroLoc can be utilized in other related tasks, like the lost-in-space problem and the localization of historical space imagery. Finally, we note that we have already used AstroLoc (in conjunction with EarthMatch’s post-processing) to provide localization for a staggering 500k photos, and anticipate that in a few months the backlog of non-localized photos will be nearly empty for the first time since the launch of the ISS.
@@ 7 Supplementary
In this Supplementary we provide additional information on the Lost-in-Space satellite problem (Sec. 7.1), on the Historical space imagery localization problem (Fig. 9), and finally we showcase a large number of qualitative results Sec. 7.3.
@@ 7.1 Lost-in-Space satellite problem
Results Tab. 4 shows AstroLoc’s superior performance in an out-of-distribution task, outperforming AnyLoc and EarthLoc by $\sim$50 R@1. These gains highlight the impact of AstroLoc’s training setup improvements. Additionally, even the smaller AstroLoc-tiny model outperforms AnyLoc with only 2% of the parameters.
Embedded Use To better understand if this model could be actually used on a nanosatellite, we also provide results with a tiny version of AstroLoc, based on the smallest version of DINO-v2 and with output dimension of 512. This version achieves 36.7% R@1 and 74.5% R@100 with only 27M parameters (25% of the full AstroLoc), making it lightweight enough to fit on an embedded system in a nanosatellite. The memory required to store database features with AstroLoc-tiny is $12k\times 512\times 4\times 4=98\text{ MB}$: 12k images with 512 dimensional features, each repeated 4 times due to the test time rotation augmentation (see Sec. 5) and each element taking 4 bytes due to float32 encoding. Further memory reduction can be achieved through compression with methods like product quantization [16] for the features or by quantizing or pruning the model itself. Given the smaller memory footprint of this model and database, it is feasible to use AstroLoc to run real time localization of nanosatellites. Combined with the state estimation techniques from  [20], AstroLoc can power a fast, accurate, and low cost orbit determination solution.
@@ 7.2 Historical space imagery localization
Historical imagery of Earth from space represents a unique source of data to understand how Earth has changed over decadal time spans. Similar to photographs taken by astronauts aboard the ISS, lots of historical imagery from space, taken from the Space Shuttle, lacks localization information. Although efforts at manually localizing these photos have been made, with over 300k photos from the early days of the Space Shuttle (1980s) manually localized, we note that (1) a large number of these images still lack location information, and (2) these images are only weakly labeled with a single location of any pixel within the image, which is often noisy and does not represent its full extent.
Given these shortcomings, we seek to understand the performance of APL systems in localizing early photographs (1981-1984) from the Space Shuttle, which were originally taken with film cameras and then digitized. As such, they have different photometric characteristics from more recent photography, and are known [^7] https://eol.jsc.nasa.gov/FAQ/#photoQuality to often require color correction (note the blue hue in many images in Fig. 9). We first precisely localize 704 images with the pipeline described in Sec. 3.1, and then compute localization results, reported in Tab. 5. We empirically find that AstroLoc achieves strong results even on these old photographs, showcasing its robustness to various types of domain changes. Comparing against performance on APL evaluation sets (Tab. 2, 3), AstroLoc has the smallest drop in recall (at all N) on the out-of-distribution historical imagery out of all methods tested.
@@ 7.3 Qualitative Results
Fig. 8, 9, 10 show qualitative AstroLoc results on astronaut photos, historical Earth from space imagery from the Space Shuttle, and satellite imagery from the McCleary[20] dataset, respectively. There are four examples in each row, with each example showing the query photo (leftmost) and the top-2 retrieval results (i.e. most similar, as ranked by AstroLoc). Correct predictions, defined as those that have any overlap with the query, are outlined in green, and incorrect predictions are outlined in red.

### v2

@@ 4.2 Query-Satellite Pairwise Loss
Given the cross-domain nature of the problem, we aim to take matching images from the different distributions and applying contrastive learning to encode their relationship in feature space. To this end, we create a batch $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{B}\}$ of query-database pairs $p_{i}=(q_{i},d_{i})\in(\mathcal{Q},\mathcal{D})$ that have an intersection over union (IoU) higher than a threshold $t_{iou}$ [^6] For the sake of readability, we abused the notation by implying that a matching query and database images from the sets $\mathcal{Q}$ and $\mathcal{D}$ have the same index as the index of the pair they belong to. This trick to simplify the notation is equivalent to have preliminarily sorted the sets $\mathcal{Q}$ and $\mathcal{D}$. (see examples in Fig. 4 top). We ensure that within a batch there are no two pairs with geographic overlap, so that for a pair $p_{i}\in\mathcal{P}$ all the images from other pairs $p_{j}\in\mathcal{P}\backslash\{p_{i}\}$ are negative examples.
@@ 4.3 Unsupervised Mining
Figure 5: Examples of training batches, using the three different sampling solutions presented. For each solution, we show two examples of batches with batch size 12 (i.e. 3 quadruplets), so that each image has 3 positives and 8 negatives. Solution 1 leads to training a non-robust model, because of the lack of hard negatives, as quadruplets are randomly sampled and are often very diverse between each other. Solution 2 provides hard negatives, but it creates batches with uninformative and feature-less images (e.g. batch 1 is from a cluster of images of seas next to coasts) which hurt the training process. Solution 3 allows to simultaneously provide hard negatives, discard uninformative batches, and focus on the most salient imagery, by sampling images from clusters that reflect the queries’ distribution.
Although the use of the Query-Satellite Pairwise Loss by itself leads to state-of-the-art results (see Sec. 5), it does not fully exploit the large quantity of available satellite imagery worldwide, as it only uses images that overlap with astronaut photographs. Therefore, we propose to use a second loss: in the next paragraphs, we explain the design of this innovative technique, starting from a naive implementation and developing into our full-fledged MUM loss.
Solution 1: naive sampling The simplest approach to train a retrieval model on large volumes of satellite imagery would be to directly apply a contrastive loss. Common training pipelines rely on using quadruplets [24, 2] (i.e. 4 images) from each class, and stacking multiple quadruplets within a batch to ensure that each image has exactly 3 positives (from the same class) and a multitude of negatives. However, this naive approach would not train a robust model, because random sampling of classes would lead to very few hard negatives (see Fig. 5 left), which are crucial to train robust image retrieval models.
Solution 2: database clustering To improve over naive sampling we could first compute features for all images within our database, and then cluster them with a k-means algorithm in feature space. This leads to the creation of $K$ clusters that share similar visual characteristics: for example one cluster containing images of forests, another one with images of deserts, and so on. These clusters can then be used to create training batches: for example there would be one batch of forest images, one batch of desert images, et cetera, leading to harder negatives within batches. While this solution leads to better results (see [6]), one issue would arise: some batches may contain only uninformative images, like feature-less images of deserts or seas, (see Fig. 5 center), which can hurt the learning process.
Solution 3: unsupervised mining We finally propose to not only create clusters from database images, but also to sample them according to the distribution of the astronaut images: intuitively this means that, as astronauts take very few photographs of deserts, and many photographs of glaciers, the clusters of desert-images will be sampled seldom, while the clusters with glaciers will be sampled more often (see Fig. 5 right).
To formalize this concept, we split the database images into a set of $K$ clusters $\mathcal{C}_{1},\ldots,\mathcal{C}_{K}$ using a standard k-means algorithm in feature space, such that
@@
Secondly, we compute features for the training queries (i.e. astronaut photos), and assign each to one of the K clusters $C$: we then denote with $b_{1},\ldots,b_{K}$ the size of the bins derived from assigning the features of the training queries to the $K$ clusters. Clusters are then sampled according to
@@
meaning that clusters with more queries are sampled more often for the creation of our training batches. We emphasize that the training batches are made only of database images, and that the queries are only used to calculate how often each cluster should be sampled, as shown in Fig. 4. Finally, given a series of quadruplets from cluster $\mathcal{C}_{K}$ (sampled according to Eq. 7), we plug them into a contrastive loss (in our case a Multi-Similarity loss [39]). Therefore, the second loss in our pipeline is
@@
where $\alpha_{2}$ and $\beta_{2}$ are positive hyperparameters, and MUM stands for Multi-similarity with Unsupervised Mining.
Among the large number of mining techniques in literature, our Unsupervised Mining stands out in two ways:
@@
    it is the only mining that uses one distribution (i.e. $Q$) to learn how to sample from another distribution (i.e. $\mathcal{D}$);
@@
    unlike common mining techniques used in localization and place recognition pipelines [3, 1, 25, 15], our miner does not require labels of queries, so we could potentially use all of the 5M existing astronaut photos (even the unlabeled ones). In practice we use only training queries to ensure avoiding any kind of test data leakage.
@@ 5.1 Implementation Details
Training We set the hyperparameters as follows: $t_{iou}=0.2,\alpha_{1}=1,\alpha_{2}=1,\beta_{1}=50,\beta_{2}=50,\lambda_{1}=1,\lambda_{2}=1,K=50$, batch size $=$ 48 (48 pairs for pair loss, 48 quadruplets for MUM loss), learning rate $=$ 5e-5, Adam optimizer. We use as architecture a DINO-v2-base backbone [28] with SALAD [17] and a linear layer to reduce feature dimensionality (from 8448 to 2048) (i.e. a similar model to AnyLoc [20] while being over 10 times lighter than AnyLoc, which uses DINO-v2-giant). Training runs for 30k iterations. As in [6], the Texas dataset is used as validation. The features for our Unsupervised Mining are computed periodically (every 5000 iterations) while the model is undergoing training, as standard practice in virtually every mining procedure [3, 1, 6, 25]. With this setup, mining increases the training time by less than 10%.
Evaluation We follow the protocol from [6] by doing image retrieval with an augmented dataset (i.e. applying four 90° rotations to each image), and using as metric the recall@N, as the percentage of queries where at least one of the top-N predictions is a correct match to the query.
@@ 5.2 Results
Results clearly show that, whereas previously there was no dominant method between EarthLoc and AnyLoc, the introduction of AstroLoc provides a new state-of-the-art model which significantly outperforms all previous, achieving a near perfect (above 99%) recall@100 on all six evaluation sets. Note that the recall@100 is more relevant in the real world than recall@1, since it is common practice to re-rank the top-N predictions with image matching methods [5, 35, 38]. Some qualitative results of queries and their predictions are in Fig. 1, with many more in the supplementary.
@@ 5.3 Other Space to Ground Use Cases
Table 4: Results of Lost-in-Space satellite problem. Using the dataset from VINSat [23] as queries, AstroLoc achieves good results on this additional task. AstroLoc-tiny is a tiny version more suitable for deployment on a nanosatellite, with 27M model parameters and 512 dimensional features (see supplementary).
@@
Historical space imagery localization Historical imagery of Earth from space is a unique source of data to understand how Earth has changed over decadal time spans. We perform experiments on 704 queries from early Space Shuttle days (1981-1984), on a worldwide database, and report results in Tab. 5. We empirically find that AstroLoc achieves strong results even on these old film photographs, showcasing its robustness to various types of domain changes.
@@ 5.4 Ablations
We compute ablations on the losses and mining in Tab. 6, where results clearly show the strong impact from each module in our pipeline. Most importantly, the combination of these components presents the optimal characteristic of orthogonality, in that their mixture has notably higher results than each of the singular elements.
Pair Loss Sol. 1 Sol. 2 Sol. 3 Texas-L Alps-L (Sec 4.2) (Sec 4.3) (Sec 4.3) (Sec 4.3) R@1 R@10 R@100 R@1 R@10 R@100 83.6 93.1 97.5 87.2 95.0 98.2 67.6 79.2 89.3 76.5 87.1 93.9 72.4 83.3 91.5 79.4 89.1 94.7 82.2 91.1 97.1 86.9 94.5 97.9 84.0 93.3 97.5 87.9 95.1 98.0 86.5 94.1 97.7 91.5 96.5 98.6 91.1 96.2 98.5 94.6 97.8 99.2
Table 6: Ablation over components and variation of the loss.
@@ 6 Conclusions
We tackle the task of Astronaut Photography Localization (APL), and through newly computed labels and ad hoc training techniques, obtain AstroLoc, which shows impressive results on all APL datasets. Furthermore, we find that AstroLoc can be utilized for other real-world applications, showing robustness to various domains. Finally, we note that we have already used AstroLoc (in conjunction with EarthMatch’s post-processing) to provide localization for a staggering 500k photos, and anticipate that in a few months the backlog of non-localized photos will be nearly empty for the first time since the launch of the ISS.
@@ 7 Supplementary
In this Supplementary we provide additional information on the Lost-in-Space satellite problem (Sec. 7.1), the Historical space imagery localization problem (Fig. 7), and finally we showcase a large number of qualitative results Sec. 7.3.
@@ 7.1 Lost-in-Space satellite problem
Results Tab. 4 (of the main paper) shows AstroLoc’s superior performance in an out-of-distribution task, outperforming AnyLoc and EarthLoc by $\sim$50 R@1. These gains highlight the impact of AstroLoc’s training setup improvements. Additionally, even the smaller AstroLoc-tiny model outperforms AnyLoc with only 2% of the parameters.
Embedded Use To better understand if this model could be actually used on a nanosatellite, we also provide results with a tiny version of AstroLoc, based on the smallest version of DINO-v2 and with output dimension of 512. This version achieves 36.7% R@1 and 74.5% R@100 with only 27M parameters (25% of the full AstroLoc), making it lightweight enough to fit on an embedded system in a nanosatellite. The memory required to store database features with AstroLoc-tiny is $12k\times 512\times 4\times 4=98\text{ MB}$: 12k images with 512 dimensional features, each repeated 4 times due to the test time rotation augmentation and each element taking 4 bytes due to float32 encoding. Further memory reduction can be achieved through compression with methods like product quantization [19] for the features or by quantizing or pruning the model itself. Given the smaller memory footprint of this model and database, it is feasible to use AstroLoc to run real time localization of nanosatellites. Combined with the state estimation techniques from  [23], AstroLoc can power a fast, accurate, and low cost orbit determination solution.
@@ 7.2 Historical space imagery localization
Historical imagery of Earth from space represents a unique source of data to understand how Earth has changed over decadal time spans. Similar to photographs taken by astronauts aboard the ISS, lots of historical imagery from space, taken from the Space Shuttle, lacks localization information. Although efforts at manually localizing these photos have been made, we note that (1) a large number of these images still lack location information, and (2) these images are only weakly labeled with a single location of any pixel within the image, which is often noisy and does not represent its full extent.
Given these shortcomings, we seek to understand the performance of APL systems in localizing early photographs (1981-1984) from the Space Shuttle, which were originally taken with film cameras and then digitized. As such, they have different photometric characteristics from more recent photography, and are known [^7] https://eol.jsc.nasa.gov/FAQ/#photoQuality to often require color correction (note the blue hue in many images in Fig. 7). We first precisely localize 704 images with the pipeline described in Sec. 3.1, and then compute localization results, reported in Tab. 5. We empirically find that AstroLoc achieves strong results even on these old photographs, showcasing its robustness to various types of domain changes. Comparing against performance on APL evaluation sets (Table 2 and Table 3 of the main paper), AstroLoc has the smallest drop in recall (at all N) on the out-of-distribution historical imagery out of all methods tested.
@@ 7.3 Qualitative Results
Fig. 6, 7, 8 show qualitative AstroLoc results on astronaut photos, historical Earth from space imagery from the Space Shuttle, and satellite imagery from the McCleary[23] dataset, respectively. There are four examples in each row, with each example showing the query photo (leftmost) and the top-2 retrieval results (i.e. most similar, as ranked by AstroLoc). Correct predictions, defined as those that have any overlap with the query, are outlined in green, and incorrect predictions are outlined in red.